if a-b=3 an a2-b2= 36, then find the sum of both the variables of a and b.
Answer by jim_thompson5910(35256)

hint: a^2 - b^2 = (a-b)(a+b)

so

a^2 - b^2 = 36

(a-b)(a+b) = 36